# An implicit–explicit BDF–Galerkin scheme of second order for the nonlinear thermistor problem

### v1

@@ 1. Introduction
Over the past decades, the thermistor problem has been widely studied in both theoretical and numerical aspects due to its significance in electrical and thermal engineering applications. Theoretical investigations have established the existence, uniqueness, and regularity properties of solutions. The existence of weak solutions was rigorously established in [Cim92] using a variational approach and techniques from functional analysis. Notably, the existence of solutions was also addressed in [AX92, AC94], while [ALM02] extended the analysis to prove the existence of weak solutions over arbitrarily large time intervals. Furthermore, the existence and uniqueness of a $C^{\alpha}$-solution in three-dimensional space were established in [Yua94].
In parallel, significant efforts have been made to develop efficient numerical methods for approximating the solution of the thermistor problem (1.1). In [EL95], optimal $L^{2}$-error estimates were derived for a semi-discrete and an implicit–explicit Euler scheme under the condition $\tau\leq\mathcal{O}(h^{d/6})$ with spatial dimension $d\leq 3$, where $\tau$ represents the time step size and $h$ denotes the mesh size. In [Zha94], a mixed finite element method (FEM) is combined with an implicit–explicit Euler scheme to obtain optimal $L^{2}$-error estimates under the time step size condition $\tau\leq\mathcal{O}(1/\sqrt{\log h^{-1}})$. Additionally, optimal $L^{2}$-error estimates were established in [AL05] for fully discrete numerical methods, where the spatial discretization was performed using a high-order FEM and the time approximation employed a combination of rational implicit and explicit multi-step schemes. However, a condition $\tau\leq\mathcal{O}(h^{3/2p})$, where $p$ is the order of time discretization, is required. In [SY18, SY19], superconvergence was obtained under the time step size restriction $\tau\leq\mathcal{O}(h^{1+\alpha})$ with $\alpha>0$. With a focus on mixed boundary conditions on Lipschitz domains, further convergence results for a wide class of finite element schemes combined with the implicit Euler scheme were established in [JMP20].
@@
Implicit–explicit methods of higher order in time only appear rarely in the literature. One positive example is [Gao16], where unconditionally optimal $L^{2}$-error estimates for a third-order implicit–explicit BDF scheme were derived. This, however, neither includes supercloseness nor superconvergence results. Building on this, we propose and analyze a fully discrete implicit–explicit BDF–Galerkin scheme for the nonlinear thermistor problem (1.1), using a variation of the second-order BDF for time discretization and standard (bi)linear finite elements for spatial discretization. The scheme is explicit in the nonlinearity and further allows a decoupled computation of the two solution components. Similar as in [AM22], this improves the computational efficiency compared to fully implicit methods. The analysis relies on the above mentioned splitting of the error components [LS13, LGS14], i.e., high-accuracy estimates for the finite element discretization and an error estimate of the numerical solution compared to the interpolation of the solution to the time-discrete system.
The remainder of the paper is organized as follows. In the forthcoming section, we introduce the necessary preliminaries and present the implicit–explicit BDF–Galerkin scheme. Sections 3 and 4 provide a detailed investigation of the error estimates in the temporal and spatial directions, respectively. Then, we collect the main results on superclose and superconvergent error estimates in Section 5. In Section 6, numerical results are presented to confirm the validity of the presented analysis. Finally, we conclude and give an outlook in Section 7.
@@ 2. Preliminaries
For an integer $k\geq 0$ and $1\leq p\leq\infty$, let $W^{k,p}(\Omega)$ denote the Sobolev space equipped with the norm $\|\cdot\|_{k,p}$ as introduced in [AF03]. For $p=2$, we write $H^{k}(\Omega)\coloneqq W^{k,2}(\Omega)$ with corresponding norm $\|\cdot\|_{k}$. Moreover, we use $(\cdot,\cdot)$ as notation for the standard $L^{2}$-inner product. To shorten notation, we further introduce the spaces
@@ 2.1. Interpolation operator and important estimates
We also make use of the subspace $V_{h,0}\coloneqq V_{h}\cap V_{0}$, which consists of discrete finite element functions that vanish on the boundary.
@@
Lemma 2.1 [Interpolation estimates, [LL07, Ch. 2]]
Suppose $u\in H^{3}(\Omega)\cap V_{0}$. Then for any $v_{h}\in V_{h,0}$ it holds that
@@
Lemma 2.2 [Inverse estimate, [BS08, Lem. 4.5.3]]
Consider parameters $1\leq q\leq p\leq\infty$ and $m\geq 0$. Then it holds that
@@
Lemma 2.3 [Discrete Gronwall inequality, [HR90]]
Let $\tau$, $B$ as well as $a_{k}$, $b_{k}$, $c_{k}$, and $\gamma_{k}$ be non-negative numbers for $k\geq 0$ such that
@@ 2.3. Implicit–explicit BDF–Galerkin scheme
Lemma 2.4 [Telescope formula for $D_{\tau}$, [Liu13]]
With the definition of the BDF operator $D_{\tau}$ in (2.10), the following telescope formula holds
@@ 3. Temporal error estimates
Lemma 3.1 [Estimate of starting approximations]
For sufficiently small $\tau$, the unique solution $(U^{1},\Phi^{1})$ of the elliptic system (3.4) satisfies the error estimate
@@
Theorem 3.1 [Temporal error estimate]
Consider $U^{0}=u^{0}$ and $(U^{1},\Phi^{1})$ from (3.4). Then, for sufficiently small $\tau$, the unique solution $(U^{n},\Phi^{n})$ of the elliptic system (3.3) satisfies the error estimate
@@
$A_{1}$ $\leq 2\,\|\sigma(U^{n-1})-\sigma(u^{n-1})\|_{0}\,\|\nabla\phi^{n}\|_{\infty}\,\|\nabla\eta_{\tau}^{n}\|_{0}$
$\leq CL_{\sigma,2}\|e_{\tau}^{n-1}\|_{0}\,\|\nabla\eta_{\tau}^{n}\|_{0}$
$\leq C\,\|\nabla e_{\tau}^{n-1}\|_{0}\,\|\nabla\eta_{\tau}^{n}\|_{0},$
@@
Using the regularity assumptions (2.9), we have together with Poincaré’s inequality
$D_{1}$ $\leq 2\,\big{(}\big{\|}\big{(}\nabla\sigma(U^{n-1})-\nabla\sigma(u^{n-1})\big{)}\cdot\nabla\phi^{n}\big{\|}_{0}+L_{\sigma,2}\,\|e_{\tau}^{n-1}\Delta\phi^{n}\|_{0}\big{)}$
$\leq 2\,\big{(}\big{(}k_{3}\|\nabla e_{\tau}^{n-1}\|_{0}+L_{\sigma^{\prime}}\|e_{\tau}^{n-1}\|_{0}\|\nabla u^{n-1}\|_{\infty}\big{)}\|\nabla\phi^{n}\|_{\infty}+L_{\sigma,2}\,\|e_{\tau}^{n-1}\|_{0}\|\Delta\phi^{n}\|_{\infty}\big{)}$
$\leq C\,\big{(}\|\nabla e_{\tau}^{n-1}\|_{0}+\|e_{\tau}^{n-1}\|_{0}\big{)}$
$\leq C\,\|\nabla e_{\tau}^{n-1}\|_{0}.$
Similarly, one obtains $D_{2}\leq C\,\|\nabla e_{\tau}^{n-2}\|_{0}$. By (3.9), it directly follows that $D_{3}\leq C\tau^{2}$. By Hölder’s inequality, we further obtain
@@
$\leq C\,\|\nabla e_{\tau}^{n-1}\|_{0,4}\|\nabla\eta_{\tau}^{n}\|_{0,4}\leq C\,\|e_{\tau}^{n-1}\|_{2}\|\nabla\eta_{\tau}^{n}\|_{2}.$
Similarly, one shows
$D_{5}\leq C\|\nabla e_{\tau}^{n-2}\|_{0,4}\|\nabla\eta_{\tau}^{n}\|_{0,4}\leq C\,\|e_{\tau}^{n-2}\|_{2}\|\nabla\eta_{\tau}^{n}\|_{2}.$
By applying elliptic regularity theory [CW98] to (3.16) and substituting the estimates for $D_{1},\ldots,D_{5}$, we obtain
@@ 4. Spatial error estimates
Lemma 4.1 [Estimates of starting approximations]
Let $(U^{1},\Phi^{1})$ denote the solution of (3.4). Then, for sufficiently small $\tau$ and $h$, the fully discrete system (2.12) with $U_{h}^{0}=I_{h}u^{0}$ has a unique solution $(U_{h}^{1},\Phi_{h}^{1})$, satisfying the error estimate
@@
Theorem 4.1 [Spatial error estimate]
Consider $(U^{n},\Phi^{n})$ as the solution of (3.3) with starting values given by (3.4). Moreover, let $(U_{h}^{n},\Phi_{h}^{n})$ denote the fully discrete solution of system (2.11) for starting values $U_{h}^{0}=I_{h}u^{0}$ and $(U_{h}^{1},\Phi_{h}^{1})$ from (2.12). Then, for sufficiently small $\tau$ and $h$, we have
@@ 5. Convergence Results
Theorem 5.1 [Convergence of the implicit–explicit BDF–Galerkin scheme]
Suppose that system (1.1) has a unique solution $(u,\phi)$ satisfying (2.9) and that $\tau$ and $h$ are sufficiently small. Then the discrete system (2.11) with the starting values $U_{h}^{1}$, $\Phi_{h}^{1}$ from (2.12) admits a unique solution $(U_{h}^{n},\Phi_{h}^{n})$, which satisfies the error bounds
(5.1a) $\|U_{h}^{n}-u^{n}\|_{0}\,+\,\|\Phi_{h}^{n}-\phi^{n}\|_{0}$ $\leq\,C\big{(}h^{2}+h\tau^{2}\big{)},$
(5.1b) $\|U_{h}^{n}-u^{n}\|_{1}\,+\,\|\Phi_{h}^{n}-\phi^{n}\|_{1}$ $\leq\,C\big{(}h+h\tau^{2}\big{)},$
(5.1c) $\|U_{h}^{n}-I_{h}u^{n}\|_{1}\,+\,\|\Phi_{h}^{n}-I_{h}\phi^{n}\|_{1}$ $\leq\,C\big{(}h^{2}+h\tau^{2}\big{)}.$
@@
Theorem 5.2 [Superconvergence]
Suppose that (2.9) is satisfied. Then, we have
$\|I_{2h}U_{h}^{n}-u^{n}\|_{1}\,+\,\|I_{2h}\Phi_{h}^{n}-\phi^{n}\|_{1}\leq C\,\big{(}h^{2}+h\tau^{2}\big{)}.$
@@ 6. Numerical results
In this section, we present numerical experiments to validate our theoretical analysis. In all computations, the final time is set to $T=1.0$. As in [LGS14, Gao16], we consider the problem
$\dot{u}-\Delta u$ $=\sigma(u)|\nabla\phi|^{2}+f_{1},$
$-\nabla\cdot(\sigma(u)\nabla\phi)$ $=f_{2}$
@@
6.1. Convergence and stability
We partition the domain $\Omega$ into squares with grid resolutions of $M\times M$ with $M=8,16,32,64$. The mesh size is given by $h=\sqrt{2}/M$. To verify the error estimates presented in Theorems 5.1 and 5.2, we first set the time step size to $\tau=\sqrt{h}$. The convergence, supercloseness, and superconvergence results for $u$ and $\phi$ at time $t=T$ are plotted in Figures 6.1 and 6.2, respectively. From Figure 6.1, we observe that $\|U_{h}^{n}-u^{n}\|_{1}$ converges at a rate of $\mathcal{O}(h)$, while $\|U_{h}^{n}-u^{n}\|_{0}$, $\|U_{h}^{n}-I_{h}u^{n}\|_{1}$, and $\|I_{2h}U_{h}^{n}-u^{n}\|_{1}$ exhibit convergence rates of order two as expected. Similarly, Figure 6.2 shows that $\|\Phi_{h}^{n}-\phi^{n}\|_{1}$ converges with order one, whereas $\|\Phi_{h}^{n}-\phi^{n}\|_{0}$, $\|\Phi_{h}^{n}-I_{h}\phi^{n}\|_{1}$, and $\|I_{2h}\Phi_{h}^{n}-\phi^{n}\|_{1}$ converge at rate $\mathcal{O}(h^{2})$. Thus, the numerical results align with the theoretical analysis of Section 5. We also present the finite element solutions $U_{h}^{n}$ and $\Phi_{h}^{n}$ in Figure 6.3 at time $t=T$ on a $32\times 32$ mesh.
[…]
Figure 6.1. Numerical errors for $u$ at final time $t=T$ using the second-order BDF–Galerkin FEM (2.11) and $\tau=\sqrt{h}$.
[…]
Figure 6.2. Numerical errors for $\phi$ at final time $t=T$ using the second-order BDF–Galerkin FEM (2.11) and $\tau=\sqrt{h}$.
Figure 6.3. Numerical solutions $U_{h}^{n}$ (left) and $\Phi_{h}^{n}$ (right) at time $t=T$ computed on a mesh with $M=32$.
To verify the unconditional convergence and accuracy of the proposed method, we conduct two complementary simulations. First, we fix the time step size $\tau$ and only refine the spatial mesh. The resulting errors for $u$ and $\phi$ (for different values of $\tau$) are displayed in Figure 6.4, demonstrating that the numerical errors stabilize as $h/\tau\rightarrow 0$ and that the proposed method does not require a time step restriction. Second, to show the temporal order of convergence, we fix the mesh size $h$, and reduce $\tau$. The $L^{2}$-errors presented in Figure 6.5 show second-order convergence, consistent with the theoretical properties of the proposed BDF–Galerkin method. Together, these results verify that the scheme achieves unconditional and optimal convergence in time and space.

### v2

@@ Abstract.
This article will be published in IMA Journal of Numerical Analysis.
@@ 1. Introduction
Over the past decades, the thermistor problem has been widely studied in both theoretical and numerical aspects due to its significance in electrical and thermal engineering applications. Theoretical investigations have established the existence, uniqueness, and regularity properties of solutions. The existence of weak solutions was rigorously established in [CIM92] using a variational approach and techniques from functional analysis. Notably, the existence of solutions was also addressed in [AX92, AC94], while [ALM02] extended the analysis to prove the existence of weak solutions over arbitrarily large time intervals. Furthermore, the existence and uniqueness of a $C^{\alpha}$-solution in three-dimensional space was established in [YUA94].
In parallel, significant efforts have been made to develop efficient numerical methods for approximating the solution of the thermistor problem (1.1). In [EL95], optimal $L^{2}$-error estimates were derived for a semi-discrete and an implicit–explicit Euler scheme under the condition $\tau\leq\mathcal{O}(h^{d/6})$ with spatial dimension $d\leq 3$, where $\tau$ represents the time step size and $h$ denotes the spatial mesh size. In [ZHA94], a mixed finite element method (FEM) is combined with an implicit–explicit Euler scheme to obtain optimal $L^{2}$-error estimates under the time step size condition $\tau\leq\mathcal{O}(1/\sqrt{\log h^{-1}})$. Additionally, optimal $L^{2}$-error estimates were established in [AL05] for fully discrete numerical methods, where the spatial discretization was performed using a high-order FEM and the time approximation employed a combination of rational implicit and explicit multi-step schemes. However, a condition $\tau\leq\mathcal{O}(h^{3/2p})$, where $p$ is the order of time discretization, is required. In [SY18, SY19], superconvergence was obtained under the time step size restriction $\tau\leq\mathcal{O}(h^{1+\alpha})$ with $\alpha>0$. With a focus on mixed boundary conditions on Lipschitz domains, further convergence results for a wide class of finite element schemes combined with the implicit Euler scheme were established in [JMP20].
@@
Implicit–explicit methods of higher order in time are very rare. To the best of our knowledge, the only positive example is [GAO16], where unconditionally optimal $L^{2}$-error estimates for a third-order implicit–explicit BDF scheme were derived. This, however, neither includes supercloseness nor superconvergence results. Building on this, we propose and analyze a fully discrete implicit–explicit BDF–Galerkin scheme of second order for the nonlinear thermistor problem (1.1), using a variation of the second-order BDF for time discretization and standard (bi)linear finite elements for spatial discretization. The scheme is explicit in the nonlinearity and further allows a decoupled computation of the two solution components. Similar as in [AM22], this improves the computational efficiency compared to fully implicit methods. The analysis relies on the above mentioned splitting of the error components [LS13, LGS14], i.e., high-accuracy estimates for the finite element discretization and an error estimate of the numerical solution compared to the interpolation of the solution to the time-discrete system. In contrast to the approach in [GAO16], we are able to prove supercloseness and superconvergence for the proposed scheme.
The remainder of the paper is organized as follows. In the forthcoming section, we introduce the necessary preliminaries and present the implicit–explicit BDF–Galerkin scheme. Sections 3 and 4 provide a detailed investigation of the error estimates in the temporal and spatial directions, respectively. Then, we collect the main results on superclose and superconvergent error estimates in Section 5. In Section 6, numerical results are presented to confirm the validity of the presented analysis and identify limitations for non-smooth solutions. Finally, we conclude and give an outlook in Section 7.
@@ 2. Preliminaries
For an integer $k\geq 0$ and $1\leq p\leq\infty$, let $W^{k,p}(\Omega)$ denote the Sobolev space equipped with the norm $\|\cdot\|_{k,p}$ as introduced in [AF03]. For $p=2$, we write $H^{k}(\Omega)\coloneqq W^{k,2}(\Omega)$ with corresponding norm
$\|\cdot\|_{k}:=\|\cdot\|_{k,2}.$
Moreover, we use $(\cdot,\cdot)$ as notation for the standard $L^{2}$-inner product. To shorten notation, we further introduce the spaces
@@ 2.1. Interpolation operator and important estimates
Note that, in this work, we only consider rectangles aligned with the coordinate axes, leading to the standard bilinear finite element space $Q_{1}$. For more general cases, isoparametric mappings are needed. We also make use of the subspace $V_{h,0}\coloneqq V_{h}\cap V_{0}$, which consists of discrete finite element functions that vanish on the boundary.
@@
Lemma 2.1
(Interpolation estimates, [LL07, Ch. 2]) Suppose $u\in H^{3}(\Omega)\cap V_{0}$. Then for any $v_{h}\in V_{h,0}$ it holds that
@@
Lemma 2.2
(Inverse estimate, [BS08, Lem. 4.5.3]) Consider parameters $1\leq q\leq p\leq\infty$ and $m\geq 0$. Then it holds that
@@
Lemma 2.3
(Discrete Gronwall inequality, [HR90]) Let $\tau$, $B$ as well as $a_{k}$, $b_{k}$, $c_{k}$, and $\gamma_{k}$ be non-negative numbers for $k\geq 0$ such that
@@ 2.2. Weak formulation and regularity assumptions
Remark 2.1.
In order to meet these strong regularity assumptions, one would expect at least a domain with $C^{2,1}$-boundary, initial data $u^{0}\in H^{3}(\Omega)$, boundary data $g\in H^{5/2}(\partial\Omega)$, as well as certain compatibility conditions between $g$ and $\phi(0)$. In the numerical experiments of Section 6, however, we will observe that second-order convergence can already be reached for less regular solutions.
@@ 2.3. Implicit–explicit BDF–Galerkin scheme
Lemma 2.4
(Telescope formula for $D_{\tau}$, [LIU13]) With the definition of the BDF operator $D_{\tau}$ in (2.10), the following telescope formula holds
@@ 3. Temporal error estimates
Lemma 3.1
(Estimate of starting approximations) For sufficiently small $\tau$, the unique solution $(U^{1},\Phi^{1})$ of the elliptic system (3.4) satisfies the error estimate
@@
Theorem 3.1
(Temporal error estimate) Consider $U^{0}=u^{0}$ and $(U^{1},\Phi^{1})$ from (3.4). Then, for sufficiently small $\tau$, the unique solution $(U^{n},\Phi^{n})$ of the elliptic system (3.3) satisfies the error estimate
@@
$A_{1}\leq 2\,\|\sigma(U^{n-1})-\sigma(u^{n-1})\|_{0}\,\|\nabla\phi^{n}\|_{\infty}\,\|\nabla\eta_{\tau}^{n}\|_{0}\leq C\,\|\nabla e_{\tau}^{n-1}\|_{0}\,\|\nabla\eta_{\tau}^{n}\|_{0},$
@@
Using the regularity assumptions (2.9), one can show with Poincaré’s inequality $D_{1}\leq C\,\|\nabla e_{\tau}^{n-1}\|_{0}$ as well as $D_{2}\leq C\,\|\nabla e_{\tau}^{n-2}\|_{0}$. By (3.9), it directly follows that $D_{3}\leq C\tau^{2}$. By Hölder’s inequality, we further obtain
@@
$\leq C\,\|\nabla e_{\tau}^{n-1}\|_{0,4}\|\nabla\eta_{\tau}^{n}\|_{0,4}\leq C\,\|e_{\tau}^{n-1}\|_{2}\|\nabla\eta_{\tau}^{n}\|_{2}$
and, similarly, $D_{5}\leq C\,\|e_{\tau}^{n-2}\|_{2}\|\nabla\eta_{\tau}^{n}\|_{2}$. By applying elliptic regularity theory [CW98] to (3.16) and substituting the estimates for $D_{1},\ldots,D_{5}$, we obtain
@@ 4. Spatial error estimates
Lemma 4.1
(Estimates of starting approximations) Let $(U^{1},\Phi^{1})$ denote the solution of (3.4). Then, for sufficiently small $\tau$ and $h$, the fully discrete system (2.12) with $U_{h}^{0}=I_{h}u^{0}$ has a unique solution $(U_{h}^{1},\Phi_{h}^{1})$, satisfying the error estimate
@@
Theorem 4.1
(Spatial error estimate) Consider $(U^{n},\Phi^{n})$ as the solution of (3.3) with starting values given by (3.4). Moreover, let $(U_{h}^{n},\Phi_{h}^{n})$ denote the fully discrete solution of system (2.11) for starting values $U_{h}^{0}=I_{h}u^{0}$ and $(U_{h}^{1},\Phi_{h}^{1})$ from (2.12). Then, for sufficiently small $\tau$ and $h$, we have
@@ 5. Convergence Results
Theorem 5.1
(Convergence of the implicit–explicit BDF–Galerkin scheme) Suppose that system (1.1) has a unique solution $(u,\phi)$ satisfying (2.9) and that $\tau$ and $h$ are sufficiently small. Then the discrete system (2.11) with the starting values $U_{h}^{1}$, $\Phi_{h}^{1}$ from (2.12) admits a unique solution $(U_{h}^{n},\Phi_{h}^{n})$, which satisfies the error bounds
(5.1a) $\|U_{h}^{n}-u^{n}\|_{0}\,+\,\|\Phi_{h}^{n}-\phi^{n}\|_{0}$ $\leq\,C\big(h^{2}+\tau^{2}\big),$
(5.1b) $\|U_{h}^{n}-u^{n}\|_{1}\,+\,\|\Phi_{h}^{n}-\phi^{n}\|_{1}$ $\leq\,C\big(h+\tau^{2}\big),$
(5.1c) $\|U_{h}^{n}-I_{h}u^{n}\|_{1}\,+\,\|\Phi_{h}^{n}-I_{h}\phi^{n}\|_{1}$ $\leq\,C\big(h^{2}+\tau^{2}\big).$
@@
Theorem 5.2
(Superconvergence) Suppose that (2.9) is satisfied. Then, we have
$\|I_{2h}U_{h}^{n}-u^{n}\|_{1}\,+\,\|I_{2h}\Phi_{h}^{n}-\phi^{n}\|_{1}\leq C\,(h^{2}+\tau^{2}).$
@@ 6. Numerical results
In this section, we present numerical experiments that validate the theoretical analysis and illustrate the performance of the proposed method. Moreover, we study a number of examples beyond the theoretical basis. In all computations, the final time is set to $T=1.0$.
6.1. Convergence and stability
In this first example, we verify the spatial and temporal convergence properties of the proposed method using a manufactured solution. To this end, artificial source terms are introduced so that the exact solution is known. As in [LGS14, GAO16], we consider the problem
(6.1a) $\dot{u}-\Delta u$ $=\sigma(u)|\nabla\phi|^{2}+f_{1},$
(6.1b) $-\nabla\cdot(\sigma(u)\nabla\phi)$ $=f_{2}$
@@ 6.1. Convergence and stability
We partition the domain $\Omega$ into squares with grid resolutions of $M\times M$ with $M=32,64,128,256$. The mesh size is given by $h=\sqrt{2}/M$. To verify the error estimates presented in Theorems 5.1 and 5.2, we first set the time step size to $\tau=h$. The convergence, supercloseness, and superconvergence results for $u$ and $\phi$ at time $t=T$ are reported in Tables 6.1 and 6.2, respectively. From Table 6.1, we observe that $\|U_{h}^{n}-u^{n}\|_{1}$ converges at a rate of $\mathcal{O}(h)$, while $\|U_{h}^{n}-u^{n}\|_{0}$, $\|U_{h}^{n}-I_{h}u^{n}\|_{1}$, and $\|I_{2h}U_{h}^{n}-u^{n}\|_{1}$ exhibit convergence rates of order two as expected. Similarly, Table 6.2 shows that $\|\Phi_{h}^{n}-\phi^{n}\|_{1}$ converges with order one, whereas $\|\Phi_{h}^{n}-\phi^{n}\|_{0}$, $\|\Phi_{h}^{n}-I_{h}\phi^{n}\|_{1}$, and $\|I_{2h}\Phi_{h}^{n}-\phi^{n}\|_{1}$ converge at rate $\mathcal{O}(h^{2})$. Note that all these numerical results align with the theoretical analysis of Section 5.
Table 6.1. Numerical errors for $u$ at final time $t=T$ using the second-order BDF–Galerkin FEM (2.11) and $\tau=h$.
mesh $32\times 32$ $64\times 64$ $128\times 128$ $256\times 256$
$\|U_{h}^{n}-u^{n}\|_{0}$ $8.33\cdot 10^{-5}$ $2.07\cdot 10^{-5}$ $5.23\cdot 10^{-6}$ $1.31\cdot 10^{-6}$
order $2.01$ $1.98$ $2.00$
$\|U_{h}^{n}-u^{n}\|_{1}$ $1.42\cdot 10^{-2}$ $7.13\cdot 10^{-3}$ $3.65\cdot 10^{-3}$ $1.82\cdot 10^{-3}$
order $1.00$ $0.97$ $1.00$
$\|U_{h}^{n}-I_{h}u^{n}\|_{1}$ $3.90\cdot 10^{-4}$ $9.52\cdot 10^{-5}$ $2.39\cdot 10^{-5}$ $5.94\cdot 10^{-6}$
order $2.03$ $1.99$ $2.01$
$\|I_{2h}U_{h}^{n}-u^{n}\|_{1}$ $2.64\cdot 10^{-3}$ $6.13\cdot 10^{-4}$ $1.50\cdot 10^{-4}$ $3.65\cdot 10^{-5}$
order $2.11$ $2.03$ $2.03$
Table 6.2. Numerical errors for $\phi$ at final time $t=T$ using the second-order BDF–Galerkin FEM (2.11) and $\tau=h$.
mesh $32\times 32$ $64\times 64$ $128\times 128$ $256\times 256$
$\|\Phi_{h}^{n}-\phi^{n}\|_{0}$ $6.75\cdot 10^{-6}$ $1.56\cdot 10^{-6}$ $3.86\cdot 10^{-7}$ $9.46\cdot 10^{-8}$
order $2.11$ $2.02$ $2.03$
$\|\Phi_{h}^{n}-\phi^{n}\|_{1}$ $1.87\cdot 10^{-2}$ $9.37\cdot 10^{-3}$ $4.71\cdot 10^{-3}$ $2.36\cdot 10^{-3}$
order $1.00$ $0.99$ $1.00$
$\|\Phi_{h}^{n}-I_{h}\phi^{n}\|_{1}$ $4.77\cdot 10^{-5}$ $1.10\cdot 10^{-5}$ $2.69\cdot 10^{-6}$ $6.58\cdot 10^{-7}$
order $2.12$ $2.03$ $2.03$
$\|I_{2h}\Phi_{h}^{n}-\phi^{n}\|_{1}$ $6.38\cdot 10^{-4}$ $1.42\cdot 10^{-4}$ $3.29\cdot 10^{-5}$ $7.93\cdot 10^{-6}$
order $2.17$ $2.11$ $2.05$
To verify the unconditional convergence and accuracy of the proposed method, we conduct two complementary simulations. First, we fix the time step size $\tau$ and only refine the spatial mesh. The resulting errors for $u$ and $\phi$ (for different values of $\tau$) are displayed in Figure 6.1, demonstrating that the numerical errors stabilize as $h/\tau\rightarrow 0$ and that the proposed method does not require a time step restriction. Second, to show the temporal order of convergence, we fix the mesh size $h$, and reduce $\tau$. The $L^{2}$-errors presented in Figure 6.2 show second-order convergence, consistent with the theoretical properties of the proposed BDF–Galerkin method. Together, these results verify that the scheme achieves unconditional and optimal convergence in space and time.